OpenStudy (anonymous):

whats the length when the area of the square is 841^2

OpenStudy (anonymous):

area of square = side^2 so side^2 = 841^2 side = 841

OpenStudy (anonymous):

The area of a square is \[a ^{2}\] so if the area is \[841^{2}\] then the side of it is also 841 (Like Harkirat just said... haha)

OpenStudy (anonymous):

lol thank you